single number from 1 to 49 take 9 steps to addition the result 50.no numbers are repeted.

Answer by AnlytcPhil(1806)

The sum of the first 9 integers is

1+2+3+4+5+6+7+8+9 = 45

We need 5 more to get 50.

If we add 10 and take out the 5 that would do the trick.

1+2+3+4+6+7+8+9+10 = 50.  That's one answer.

Or we could add ll and take out the 6

1+2+3+4+5+7+8+9+11 = 50.  That's another answer.

Or we could add l2 and take out the 7

1+2+3+4+5+6+8+9+12 = 50.  That's another answer.

We can make another one out of that one, by replacing
the 9+12 by 10+11

1+2+3+4+5+6+8+10+11 = 50.  That's another answer.

Or we could add l3 and take out the 8

1+2+3+4+5+6+7+9+13 = 50.  That's another answer.

We can make another one out of that one, by replacing
the 9+13 by 10+12

1+2+3+4+5+6+7+10+12 = 50.  That's another answer.

Or we could add l4 and take out the 9

1+2+3+4+5+6+7+8+14 = 50.  That's another answer.

So that gives you 7 different answers to get 50.

Edwin